PBFIFO(8) Linux PBFIFO(8) NAME pfifo−PacketlimitedFirstIn,FirstOutqueue bfifo−BytelimitedFirstIn,FirstOutqueue SYNOPSIS tcqdisc...addpfifo[limitpackets] tcqdisc...addbfifo[limitbytes] DESCRIPTION ThepfifoandbfifoqdiscsareunadornedFirstIn,FirstOutqueues.Theyarethesimplestqueues possibleandthereforehavenooverhead.pfifoconstrainsthequeuesizeasmeasuredinpackets. bfifodoessoasmeasuredinbytes. Likeallnon-defaultqdiscs,theymaintainstatistics.Thismightbeareasontopreferpfifoorbfifo overthedefault. ALGORITHM Alistofpacketsismaintained,whenapacketisenqueueditgetsinsertedatthetailofalist. Whenapacketneedstobesentouttothenetwork,itistakenfromtheheadofthelist. Ifthelististoolong,nofurtherpacketsareallowedon.Thisiscalled’taildrop’. PARAMETERS limitMaximumqueuesize.Specifiedinbytesforbfifo,inpacketsforpfifo.Forpfifo,defaultsto theinterfacetxqueuelen,asspecifiedwithifconfig(8)orip(8). Forbfifo,itdefaultstothetxqueuelenmultipliedbytheinterfaceMTU. OUTPUT Theoutputoftc-sqdisclscontainsthelimit,eitherinpacketsorinbytes,andthenumberof bytesandpacketsactuallysent.Anunsentanddroppedpacketonlyappearsbetweenbracesand isnotcountedas’Sent’. Inthisexample,thequeuelengthis100packets,45894bytesweresentover681packets.No packetsweredropped,andasthepfifoqueuedoesnotslowdownpackets,therewerealsonooverlimits: #tc-sqdisclsdeveth0 qdiscpfifo8001:deveth0limit100p Sent45894bytes681pkts(dropped0,overlimits0) Ifabacklogoccurs,thisisdisplayedaswell. SEEALSO tc(8) AUTHORS AlexeyN.Kuznetsov,<kuznet@ms2.inr.ac.ru> Thismanpagemaintainedbyberthubert<ahu@ds9a.nl> iproute2 10January2002 1